OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7).

OpenStudy (anonymous):

@d3Xter

OpenStudy (anonymous):

Given the linear equation 3x + y = 5, find the slope of its graph. A. 3 B. –5 C. –3 D. 5

OpenStudy (anonymous):

arrange your equation in the point-slope eqn form y = mx + b Now what is m from your result?

OpenStudy (anonymous):

are you talking about the first question or the second question

OpenStudy (anonymous):

the second

OpenStudy (anonymous):

well when I first did it I Thought maybe it was -3

OpenStudy (anonymous):

And, the correct expression is slope-intercept eqn. sorry

OpenStudy (anonymous):

yeah you're right! let's rearrange that y = -3x + 5 so m = -3 = slope

OpenStudy (anonymous):

and b = 5 = intercept

OpenStudy (anonymous):

so what about the first one

OpenStudy (anonymous):

we still making use of the slope-intercept eqn here, but first we'll find the slope m\[m = \frac{ rise }{ run } = \frac{ y - y_{1} }{ x - x _{1}}\]

OpenStudy (anonymous):

Are you familiar with that equation?

OpenStudy (anonymous):

ya I think its y=2/3x+13/3

OpenStudy (anonymous):

hold on, let me check that.

OpenStudy (anonymous):

I think it is at least similar to that

OpenStudy (anonymous):

\[m = \frac{ 7 - 1}{ -4 -5} = -\frac{ 6 }{ 9 } = -\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

then our eqn becomes \[y = -\frac{ 2 }{ 3 }x + b\]

OpenStudy (anonymous):

is it possible b=13/3

OpenStudy (anonymous):

using the equation y = y1 + m(x - x1) let's see! we'll use the first coordinate (5, 1) -> (x1,y1) \[y = 1 - \frac{ 2 }{ 3 }(x -5) = 1 - \frac{ 2 }{ 3 }x + \frac{ 10 }{ 3 } \rightarrow -\frac{ 2 }{ 3 }x + \frac{ 13 }{ 3 }\]

OpenStudy (anonymous):

So, yeah. We're right! :)